OpenStudy (anonymous):

multiply x^1/4 times x^5/8

OpenStudy (anonymous):

\[x^\frac{ 1 }{ 4 }\times x^\frac{ 5 }{ 8 }\]

OpenStudy (anonymous):

i think i have to put it in radical form

OpenStudy (anonymous):

but i dont know what to do from there

OpenStudy (anonymous):

add the exponents

OpenStudy (anonymous):

so the exponents would be 7 right?

OpenStudy (anonymous):

$$\frac{1}{4}+\frac{5}{8}$$

OpenStudy (anonymous):

so 1/2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no it would be $$\frac{7}{8}$$